OpenStudy (henrietepurina):

The equation below represents Function A and the graph represents Function B. Function A f(x) = x - 9 Function B

OpenStudy (henrietepurina):

Which equation best compares the slopes of the two functions? Slope of Function B = 2 × Slope of Function A. Slope of Function A = Slope of Function B Slope of Function A = 2 × Slope of Function B Slope of Function B = - Slope of Function A

OpenStudy (kc_kennylau):

Okay, find the slope of function B first

OpenStudy (kc_kennylau):

Slope is rise over run, or y-difference divided by x-difference

OpenStudy (kc_kennylau):

Use any two points on the line to find the slope

OpenStudy (henrietepurina):

ok so...

OpenStudy (henrietepurina):

how if there re no points

OpenStudy (kc_kennylau):

You create two points on the line

OpenStudy (kc_kennylau):

(0,-1) and (1,1) are good points

OpenStudy (henrietepurina):

ok so -1 - 1 = m(0 - 1)

OpenStudy (henrietepurina):

-2 = m(-1)

OpenStudy (henrietepurina):

-2 = -m

OpenStudy (kc_kennylau):

And m=?

OpenStudy (henrietepurina):

2?

OpenStudy (kc_kennylau):

Yep

OpenStudy (kc_kennylau):

Find the slope of function A?

OpenStudy (henrietepurina):

-7?

OpenStudy (kc_kennylau):

Why?

OpenStudy (henrietepurina):

no i meant -9

OpenStudy (henrietepurina):

or wait -5

OpenStudy (kc_kennylau):

Why and why?

OpenStudy (henrietepurina):

well bec -2 - 9

OpenStudy (henrietepurina):

equals... ohhhh, 7?

OpenStudy (kc_kennylau):

Why -2-9?

OpenStudy (kc_kennylau):

Remember that in the form y=mx+b, m is the slope and b is the intercept

OpenStudy (henrietepurina):

omg its getting late just cut to the poop

OpenStudy (kc_kennylau):

So in f(x)=x-9, the slope is 1

OpenStudy (kc_kennylau):

So the answer is A

OpenStudy (henrietepurina):

which one a b c or d

OpenStudy (henrietepurina):

ok thanks